### v2

# thanks: M. A. C. A., N.-Y. L. and C.-H. M. contributed equally in the manuscriptthanks: M. A. C. A., N.-Y. L. and C.-H. M. contributed equally in the manuscriptthanks: M. A. C. A., N.-Y. L. and C.-H. M. contributed equally in the manuscript

Abstract
Superconducting qubits have achieved exceptional gate fidelities, exceeding the error-correction threshold in recent years. One key ingredient of such improvement is the introduction of tunable couplers to control the qubit-to-qubit coupling through frequency tuning. Moving toward fault-tolerant quantum computation, increasing the number of physical qubits is another step toward effective error correction codes. Under a multiqubit architecture, flux control (Z) lines are crucial in tuning the frequency of the qubits and couplers. However, dense flux lines result in magnetic flux crosstalk, wherein magnetic flux applied to one element inadvertently affects neighboring qubits or couplers. This crosstalk obscures the idle frequency of the qubit when flux bias is applied, which degrades gate performance and calibration accuracy. In this study, we characterize flux crosstalk and suppress it in a multiqubit-coupler chip with multi-Z lines without adding additional readout for couplers. By quantifying the mutual flux-induced frequency shifts of qubits and couplers, we construct a cancellation matrix that enables precise compensation of non-local flux, demonstrating a substantial reduction in Z-line crosstalk from 56.5 ‰ to 0.13 ‰  which is close to statistical error. Flux compensation corrects the CZ SWAP measurement, leading to a symmetric map with respect to flux bias. Compared with a crosstalk-free calculated CZ SWAP map, the measured map indicates that our approach provides a near-zero crosstalk for the coupler-transmon system. These results highlight the effectiveness of our approach in enhancing flux crosstalk-free control and supporting its potential for scaling superconducting quantum processors.
Keywords: Superconducting qubits, Quantum Physics, Quantum Information
preprint: APL Quantum
Superconducting qubits are one of the promising types of physical qubits to realize fault-tolerant quantum computing (FTQC), in which quantum computers tackle problems that conventional computers cannot solve on a reasonable time scale [1, 2]. In operating high-fidelity quantum gates, the tunable coupler plays a crucial role. It suppresses spurious coupling and modulates the coupling strength between two qubits [3, 4, 5] via coupler-mediated virtual exchange [4, 1]. These features enhance single-qubit (1Q) and two-qubit (2Q) gate fidelities in both transmon and Fluxonium architectures  [6, 7] in a superconducting quantum processor. Reaching the error correction threshold marks the first step towards FTQC  [1, 8, 9, 10].
To achieve FTQC, an increasing number of qubits and couplers is needed, which leads to more control lines as well as crosstalk between qubits and couplers. Crosstalks could originate from unintended microwave [11, 12, 13, 14] and readout drives [15, 16], residual qubit-to-qubit coupling [3, 4, 6, 17], frequency crowding [18, 5, 19, 20], and unwanted flux [21, 22, 23, 24] that leads to imperfection of the 1Q and 2Q gates. Flux crosstalk plays a role in a frequency-tunable qubit-coupler system; it obscures the precise frequency detuning required for a quantum gate operation. Precise control of the flux pulse avoids unwanted frequency shifts and residual coupling, which is beneficial for quantum gate operation.
As flux-tunable elements are sources of flux crosstalk in a superconducting processor, the crosstalk matrix shows the spatial effect of the flux source based on the frequency response of other probed elements [23, 25, 26]. The diagonal elements denote the flux source, whereas the off-diagonal elements represent the flux crosstalk. Frequency-sensitive techniques like qubit spectroscopy [23, 27], resonator spectroscopy [28, 29, 24], and Ramsey sequences [22, 6, 30] all characterize DC components. Phase-sensitive techniques such as Ramsey sequences [28, 27] capture high-frequency components. Adding compensating pulses based on the inverse matrix suppresses these crosstalk [21, 22]. Flux crosstalk compensation, along with flux pulse predistortion [6, 17, 31], enables the prediction of idle frequencies for qubits and couplers in large quantum processors [21, 25, 31].
In a transmon-coupler (TC) architecture [4, 6, 1, 5], flux control relies on the current flowing through a flux (Z) line grounded near the qubit or coupler. As depicted in Fig. 1(a), four flux lines, colored green and labeled Z, are responsible for frequency control with flux threading through the superconducting quantum interference device (SQUID) loop. To characterize flux crosstalk in the TC system, we utilize the tunable two-qubit and two-coupler system, as shown in Fig. 1(a). Here, the element referring to the flux line $i,k\in\{\textrm{{$Z_{\textrm{Q1}}$},{$Z_{\textrm{Q2}}$}\,$Z_{\textrm{C1}}$},\textrm{$Z_{\textrm{C2}}$}\}$ interacts with an element referring to the SQUID loop of a target qubit or coupler $j\in\{\textrm{Q1},\textrm{Q2},\textrm{C1},\textrm{C2}\}$. When the current $I_{i}$ generated by a voltage difference $V_{i}$ across a load impedance of $50\,\Omega$ passes through a target flux line antenna, $I_{i}$ induces two effects that generate flux crosstalk: 1) residual inductive coupling $L_{ij,i\neq j}$ between flux lines and SQUID loops of neighboring qubits and couplers and 2) unwanted fluxes threading through the SQUID structure due to return currents [23, 27]. Flux crosstalk complicates the precise control of TC system for quantum gate operations.
To overcome the crosstalk challenge, we develop a protocol for characterizing flux crosstalk in a compact TC architecture that exploits residual inductive coupling between several flux-tunable elements to resolve the 0–1 transition frequencies of the probed element, named ”Multi-Z-Line Control” (MZLC). The method only requires prior knowledge of the flux bias dependence of the qubit/coupler spectrum. It works even with coarsely optimized readout fidelity and applies to arbitrary flux biases of the probed element. We apply this method in compact TC with dense wiring of flux lines without a dedicated readout resonator and drive line for coupler. MZLC reveals average flux crosstalk of $27\,\pm\,19\,$‰  and $0.2\,\pm\,0.1\,$‰, before and after the flux crosstalk matrix compensation. Flux crosstalk matrix cancellation after MZLC enables the coupler-assisted coupling strength to be symmetric with varying coupler detuning after compensation, easing up the tune-up of conditional-phase (CZ) gates.
Figure 1: Subsystem used to characterize magnetic flux crosstalk. (a) Illustration of flux crosstalk. Magnetic flux from the Z-lines causes unwanted frequency shifts of nearby tunable elements. (b) Optical micrograph of the chip, which has five flux-tunable transmon qubits with readout resonators and four flux-tunable couplers. (c) Optical micrograph of the subsystem consisting of two qubits (Q1, Q2) in blue and two couplers (C1, C2) in orange. The inset shows a zoomed-in picture of the device with visual termination of the flux lines adjacent to the SQUID. Color labels are assigned based on the corresponding components in the subsystem’s lumped circuit model shown in (d).
The protocol for characterizing flux crosstalk is carried out on a subsystem of a quantum processing unit (QPU) containing five qubits and four couplers, as shown in Figs. 1(b). The QPU is fabricated with conventional e-beam lithography, followed by a tilt-angle deposition of aluminum with in situ oxidation to make the circuit elements and Manhattan-style Josephson junctions. The MZLC protocol is implemented in a TC subsystem located at the edge of the linear qubit chain, as shown in Figs. 1(b) and 1(c). Wire bonds are placed above the microwave and flux lines to suppress slot line modes [32], and flux crosstalk [27]. The experiment is conducted in a dilution refrigerator with a base temperature of $10\text{\,}\mathrm{m}\mathrm{K}$, and the microwave signal is supplied by Quantum Machines OPX++\, along with Octave up/down converters (see Sec. SII of [33] for additional details).
[…]
Next, we detail the protocol for characterizing the crosstalk of the TC system. Figures 2(a) and 2(b) show standard two-tone spectroscopy pulse schedules for qubits and couplers, respectively, with the same color codes as Figs. 1. Figures 2(c) and 2(e) reveal the frequency spectrum of the qubits Q1 and Q2, with the driving frequency $f_{\text{XY}}$ as a function of the flux bias voltage $V_{i}$. Note that every element involved in the pulse sequences shown in all figures may subject to a constant DC offset component of $V_{i}$ to compensate for the flux offset induced by trapped magnetic fields during cooldown. Any variations in the flux bias voltage amplitude shown in all pulse sequences correspond to the pulsed flux component of $V_{i}$.
Due to the absence of dedicated readout resonators and XY lines for the couplers, we employ indirect coupler spectroscopy (ICS) with the pulse sequence shown in Fig. 2(b). ICS exploits the weak capacitive coupling between the coupler electrode and the XY and readout lines of the adjacent qubits to resolve $f_{01,\text{C}}$ [34, 35]. In Fig. 2(b), the XY lines of Q1 and Q2 drive the 0–1 transition of C1 and C2, respectively. Then R1 and R2 read the 0–1 transition of C1 and C2, respectively. Figs. 2(d) and 2(f) show the frequency spectra of couplers C1 and C2, respectively. In obtaining Figs. 2(d) and 2(f), the coupler is driven with elevated amplitude of microwave and readout tones to better resolve the 0–1 transition spectrum. Moreover, adding appropriate delay to overlapping drive and readout pulses enhances the contrast of the 0–1 transition in the coupler spectrum, thereby improving the signal-to-noise ratio for both qubit spectroscopy and ICS. As these weak capacitive couplings are inherent in the subsystem design, ICS is compatible with standard measurement schemes.
Sweeping the microwave drive frequency $f_{\text{XY}}$ of the qubits’ XY line with fixed XY drive and flux bias reveals a peak frequency response with a Lorentzian profile of both the qubits and couplers. The peak frequency, representing the 0–1 transition, changes with flux bias voltage $V_{i}$, and follows $f_{\text{01}}$ of the flux-tunable transmon model, which is shown as green dashed lines in Figs. 2(c) to 2(f). The model is written as  [36, 23]
$\begin{split}f_{01}(V_{i})&=\left(f_{01,\max}+\frac{E_{C}}{h}\right)\times\\ &\quad\sqrt[4]{d^{2}+\left[1-d^{2}\right]\textrm{cos}^{2}\left(A_{c}\left[V_{i}-V_{\text{ofs,i}}\right]\right)}\\ &\quad-\frac{E_{C}}{h},\end{split}$ (1)
where $f_{\text{01,max}}$ is the upper sweet spot frequency, $E_{C}/h$ is the anharmonicity, $d$ is the junction asymmetry, $A_{c}$ is a conversion factor related to the mutual inductance between the $i^{th}$ SQUID loop and the $i^{th}$ flux line, and $V_{\text{ofs,}i}$ is the offset voltage induced by the remnant flux. Notably, $A_{c}$ and $V_{\text{ofs,}i}$ are related to the normalized flux of the transmon by $A_{c}(V_{i}-V_{\text{ofs,}i})=\pi\Phi/\Phi_{0}$ where $\Phi$ is the total flux that traverses the SQUID loop and $\Phi_{0}$ is the flux quanta. The transmon anharmonicity is determined by two-photon excitation measurements. The junction asymmetries for both qubits and couplers are based on the nominal design in Sec. SI of [33]. Hence, $A_{c}$, $f_{\text{01,max}}$ and $V_{\text{ofs,}i}$ are extracted when fitting the model with the driven spectra in Figs. 2(c) and 2(f), expressed in voltages from the analog-to-digital converter in-phase channel (ADC-I) [37]. These parameters map the idle transmon frequency to its corresponding flux bias  [36], and vice versa after flux crosstalk compensation (see Sec. SIII A in [33] for details).
Figure 3: Characterization of flux crosstalk. (a) Pulse sequence used to extract crosstalk of MZLC: a probing element is driven at a fixed frequency, while both probe voltage and source voltage are applied simultaneously. The XY, Z, and readout pulses have $20\text{\,}\mathrm{n}\mathrm{s}$ delays to avoid timing mismatch artifacts. Spectra with crosstalk are illustrated in intensity plots with C1 as source and Q1 as probe (b) and vice versa (d). The horizontal, vertical, and intensity axes are the source flux (S-Flux), probe flux (P-Flux), and readout signal voltage, respectively. Applying flux-compensated pulses to (b) ($0.39\times\text{P-Flux}_{\text{Q1}}$) and (d) ($0.34\times\text{P-Flux}_{\text{C1}}$) after characterizing the flux crosstalk results in spectra immune to S-Flux as shown in (c) ($0.39\times\text{P-Flux}_{\text{C1}}$) and (e) ($0.33\times\text{P-Flux}_{\text{Q1}}$), respectively. In (f) ($0.40\times\text{P-Flux}_{\text{C1}}$) and (g) ( $0.24\times\text{P-Flux}_{\text{C1}}$), as the frequency detuning of probe C1 from its sweet spot frequency of $f_{\text{C1,max}}(\Phi_{p}=0)=8046.6\,\pm\,0.2\,$MHz decreases, the spectral linewidth of C1 increases. All measurement conditions in this figure have an identical readout pulse buffer and initialization time.
We then investigate the flux crosstalk for the qubits and couplers. Figures 3(a) shows the pulse sequence of the MZLC protocol with a similar legend in Fig. 2(a). Before the protocol begins, the XY pulses for both the probe and source elements are calibrated at the maximum transmon frequency. Next, the DC flux biases of the probe and source are set to their respective target operating points. The XY pulse frequency is then fixed to match the 0–-1 transition at the target operating point. To characterize the flux crosstalk, we simultaneously sweep the probe flux pulse amplitude (P-Flux) and the source flux pulse amplitude (S-Flux), ensuring that both pulses have the same duration. To prevent timing mismatch artifacts, a $20\text{\,}\mathrm{n}\mathrm{s}$ buffer time is inserted between the flux, XY, and readout pulses. After the XY pulse is turned off, an additional 20 ns buffer is applied before both the S-Flux and P-Flux pulses are diabatically turned off. A $20\text{\,}\mathrm{n}\mathrm{s}$ delay is also introduced between the flux and readout pulses to reduce transient effects from the flux pulses. Finally, a readout pulse from the probe is used to resolve the probe frequency.
Figure 4: Flux crosstalk matrices between tunable elements. Rows correspond to detector elements, and columns to source Z-lines. (a) Displaying results obtained from two-tone spectroscopy averaged over 100 repetitions to estimate statistical uncertainty. (b) shows corresponding results from Ramsey, which agrees with statistical error to (a). (c) show the crosstalk matrix after compensation, demonstrating suppression of all off-diagonal elements to below 0.50.5\,‰. Values smaller than their uncertainties are kept to show the nominal measurement mean, which otherwise rounds to zero.
[…]
Figures 3(b) and 3(d) show crosstalk measurement results using MZLC. The horizontal and vertical axes represent the applied voltage on the source and probe flux lines, respectively, which are converted to flux quanta, named S-Flux ($\pi\Phi_{s,i}/\Phi_{0}$) and P-flux ($\pi\Phi_{p,i}/\Phi_{0}$). Each data point represents a frequency response ($f_{\text{01}}$) of the qubits (couplers) as a detector under a fixed XY frequency, driven at the nominal qubit frequency extrapolated from the fit in Figs. 2(c) and 2(d), and a simultaneous bias of the source and probe elements.
MZLC resolves the behavior of the 0–1 transition of the probe by monitoring changes in its population, which is between $|0\rangle$ and $|1\rangle$ for a weak pulsed XY drive, with different combinations of P-Flux and S-Flux. The combined pulses will allow additional sampling of the flux crosstalk effect with better precision, and can ensure the linear correlation between probe and source elements, instead of 1D scans as done in other works [22]. The scheme hinges on detecting flux crosstalk by observing undesirable frequency shifts on the detector induced by S-Flux, which is visually distinct from other sources of unwanted frequency shifts. In the absence of flux crosstalk, the 0–1 transition of the probe has a characteristic P-Flux that is unaffected by the S-Flux sweep. With flux crosstalk, the 0–1 transition appears to vary linearly with the S-Flux sweep. The magnitude of the slope captures the total applied flux magnitudes as done previously with other techniques [22]. Other sources of frequency shift, such as dispersive shifts caused by neighboring flux-tunable elements, and ZZ interaction-induced shift, appear as a nonlinear dependence of the probe’s 0–1 transition [23] at different combinations of simultaneously biased P-Flux and S-Flux.
It is worth noting that an alternative method is the Ramsey-based method [22, 6, 30], which observes the same sign and magnitude of crosstalk, as shown in Fig. 4(b) and Fig. S2 of the supplemental material [33]. Within the S-flux ranges shown in Figs. 3(b) and 3(d), the $f_{\text{01}}$ of the probe elements are linearly proportional to the flux quanta of the source element, implying linear crosstalk between neighboring flux-controlled elements. We model this linear dependence on the source frequency $f_{s,i}$ as
$f_{p,k}=X_{ik}f_{s,i}+f_{p,i}=\frac{-\Delta V_{p,k}}{\Delta V_{s,i}}f_{s,i}+f_{p,i},$ (2)
where $X_{ik}=-\Delta V_{p,k}/\Delta V_{s,i}$ is the flux crosstalk coefficient obtained from the slope of Figs. 3(b) and 3(d). We label $X_{ik}$ as the non-diagonal elements of the flux crosstalk matrix $X$ along with its normalized diagonal term ($X_{ii}=1$). With $X$ measured, one can calculate the inverse matrix, which leads to a cancellation matrix $X^{-1}$. After applying $X^{-1}$ to reconfigure our flux pulses, the remeasured flux crosstalk maps in Figs. 3(c) and 3(e) show that the S-Flux does not affect on the transition frequency of the probe element. The results not only support the validity of Eq. (2) but also confirm a significant suppression of flux crosstalk. Notably, the intercept $f_{p,i}$ extracted from Figs. 3(b) and 3(d) is a good estimate of the probe idle frequency $f_{p,k}$ after flux compensation, as shown in Figs. 3(c) and 3(e). We attribute these megahertz discrepancies to uncorrected long-time-scale transients of the flux pulses [6, 31, 17] (see Fig. S4 and Sec. SIII F of [33] for details).
These megahertz discrepancies, if not corrected first with flux predistortion and then flux crosstalk compensation, lead to coherent errors that affect the 1Q gate fidelities of qubits in a multi-qubit processor having simultaneous flux pulse biasing (see Secs. SVI and SVII A of [33] for details).
When using the MZLC approach, detuning the probe element away from its upper sweet spot leads to a narrow linewidth, which improves the precision of flux crosstalk measurement, as shown in Fig. 3(f). Conversely, detuning the probe element towards the upper sweet spot frequencies, as shown in Fig. 3(g), results in a broadened spectral linewidth and larger error bar in the crosstalk fitting. These linewidths are consistent with the qubit and coupler spectra in Fig. 2, which are observed in previous works [38, 39, 40, 6, 41].
In fast-flux qubit spectroscopy, whose pulse time is much faster than the relaxation and dephasing time $t\ll T_{1},T_{2}^{*}$, the linewidth of the qubit spectrum is Fourier-limited by the inverse of the pulse duration time. However, the XY Rabi drive, which is proportional to the transmon’s dipole matrix element, and consequently the fourth root of the Josephson energy [42], has periodic dependence on the applied flux. The flux-dependent XY Rabi drive modulates the effective qubit linewidth. As our detection undergoes repeated ensemble averaging per pixel to resolve the 0–1 transition, low-frequency drifts (like qubit frequency drifts) are averaged out, leading to qubit dephasing [43]. At a weak XY drive, the observed pulsed spectroscopy lineshape becomes Gaussian-like due to inhomogeneous broadening [44, 43, 45] (see Sec. SIV of [33] for details). Despite the observed broadening, peak thresholding and curve fitting improve the frequency shift resolution (see Sec. SV of [33] for details). Nonetheless, the sign and amplitude of the crosstalk in Figs. 3(f) and 3(g) remain unchanged.
Table 1: Metrics for Flux Crosstalk $X_{ik}$ (in ‰) before and after flux crosstalk compensation based on Figs. 4(a) and 4(c).
Metric Notation Before After
Largest Negative $\text{Min}(X_{ik})$ $58\,\pm\,1$ $0.5\,\pm\,0.4$
Largest Positive $\text{Max}(X_{ik})$ $43\,\pm\,2$ $0\,\pm\,1$
Average $\langle|X_{ik}|\rangle_{i\neq k}$ $27\,\pm\,19$ $0.2\,\pm\,0.1$
Total $\sum_{i\neq k}|X_{ik}|$ 318.4 2.42
Matrix
Asymmetry
$\sum_{i\neq k}$
$||X_{ik}|-|X_{ki}||$
143.6 0.94
We summarize the crosstalk matrices measured using the MZLC and Ramsey sequences in Figs. 4(a) and 4(b), respectively. Both methods yield consistent results within the statistical error. We discuss the efficiency of MZLC and Ramsey schemes in Secs. SIII B, and SVII of [33]. Table 1 shows the statistical properties of $X_{ik}$ before and after flux compensation in Figs. 4(a) and 4(c), respectively. Few observations of the crosstalk matrix are as follows. First, the average uncompensated crosstalk has the same order as its extreme values (Fig. S3 of [33] presents the same results in dB). Its total crosstalk is similar to previous flux-tunable two-qubit transmons [46]. Its matrix asymmetry exhibits an imbalance likely stemming from the leftward routing layout of our flux lines, producing asymmetric inductive coupling and return current. A careful routing of flux lines [27] with the return current flowing directly on the PCB ground [47] could passively reduce the flux crosstalk and matrix asymmetry to a certain extent. Lastly, these metrics dropped one to two orders of magnitude after flux compensation; residual crosstalk averages near zero, with a standard deviation of a similar scale. The flux-compensated crosstalk becomes decoupled, uniform, and reciprocal.
Figure 5: Time-domain measurements of the coherent oscillations between $|11\rangle$ (red) and $|02\rangle$ (blue) states, named CZ SWAP, under varying coupler flux and different flux pulse durations. The pulse schedule in (a), with the XY drive being the $\pi_{01}$ pulse, and the results before flux compensation (b) and after flux compensation (c) are displayed. The flux-compensated time-domain measurements agree with simulations in (d). (e) 2D Fast Fourier transform (FFT) of the time-domain oscillations in (b), (c), and (d), which quantifies the conditional phase-mediated coupling strength $|g_{\mathrm{eff}}|$ as a function of coupler flux. Inset shows the residuals (dotted lines) expressed in MHz. Panels from (b) to (e) share the same x-axis, representing the applied flux on the coupler.
In addition to compensating for flux crosstalk, we applied this method to help locate the optimal bias point for conditional phase (CZ) gate operations [48, 6, 17]. In TC architectures, the coupler-mediated CZ gate works by tuning the frequencies of the qubits so that the states $|11\rangle$ and $|20\rangle$ become resonant, resulting in a state swap. The effective coupling strength $g_{\mathrm{eff}}$ in this regime is susceptible to detuning between the frequencies of qubits and couplers, which in turn depends on the applied flux bias.
We show measured coupler-tuned CZ SWAP experiments using the pulse schedule as shown in Fig. 5(a) (see Sec. SIII D of [33] for the relevant CZ SWAP idle frequencies). The behavior seen in Figs. 5(b) and 5(c) before and after flux compensation, respectively, are compared with Fig. 5(d), which is derived from the Rabi oscillation for 2Q gates [41, 49], accounting for the finite rising time of the diabatic flux pulse. This finite rise time—an effective flux transient representing the combined, rather than individual, components of flux pulse distortions—results in fewer Rabi oscillation cycles and, consequently, a reduced interaction time for two-qubit (2Q) gates (see additional details of the model from Secs. SIII D to SIII F of [33]). Due to phase errors arising from the uncalibrated pulse distortions in flux biasing [25, 17], Figs. 5(b) and 5(c) may show a small deviation from Fig. 5(d).
[…]
The CZ coupling strength in a TC architecture (Q1-C2-Q2) is written as [50, 51]
$g_{\text{eff}}=\sqrt{2}\left(g_{12}-\frac{g_{1c}g_{2c}}{2}B_{02}\right),$ (3)
where $g_{12}$ is the direct coupling between Q1 and Q2, $g_{1c}$ is the coupling strength between Q1 and coupler C2, $g_{2c}$ is the coupling strength between Q2 and coupler C2, and
$B_{02}=\frac{1}{\Delta_{1}}+\frac{1}{\Delta_{2}-E_{C,\text{Q2}}/\hbar}+\frac{1}{\sum_{1}}+\frac{1}{\sum_{2}+E_{C,\text{Q2}}/\hbar},$ (4)
where $\Delta_{1}/2\pi=f_{01,\text{C2}}-f_{01,\text{Q1}}$ and $\Delta_{2}/2\pi=f_{01,\text{C2}}-f_{01,\text{Q2}}$ are the frequency detunings between the coupler and the qubits, and $\sum_{1}/2\pi=f_{01,\text{C2}}+f_{01,\text{Q1}}$ and $\sum_{2}/2\pi=f_{01,\text{C2}}+f_{01,\text{Q2}}$ are the frequency sums of the transition frequencies. Moreover, $E_{C,\text{Q1}}/\hbar$ and $E_{C,\text{Q2}}/\hbar$ are the anharmonicities of Q1 and Q2, respectively. To fit the flux-compensated $|g_{\text{eff}}|$ in Fig. 5(e), we use Eqs. (3)–(4) with an added coupling strength of |435||435|\,kHz, which could be reduced through better readout fidelity and rigorous flux predistortions [17, 25, 49]. The inset in Fig. 5(e) accentuates the effect of flux compensation on predicting $|g_{\text{eff}}|$. At $|\Phi_{\text{C2}}|<0.88\Phi_{0}$, uncompensated flux pulses generate asymmetric $|g_{\text{eff}}|$ with megahertz offsets from the fit, whereas compensated flux pulses produce predictable $|g_{\text{eff}}|$.
Non-idealities in Fig. 5(e) and its inset become evident when comparing the measured $|g_{\text{eff}}|$ with the fit at $|\Phi_{\text{C2}}|\geq 0.88\Phi_{0}$. The fit cannot explain the sharp peaks at $|g_{\text{eff}}/2\pi|$ at $|\Phi_{\text{C2}}|\approx 0.88\Phi_{0}$ ($f_{\text{C2}}\approx$ $5.94\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$) as the predicted singularities occur beyond the range of the coupler flux displayed in Fig. 5(e) [51]. The singularities probably come from the strong coupling between the readout resonator R2 and the coupler C2 as $f_{\text{C2}}\approx f_{\text{R2}}$ (see Table S2 of [33]), which produces hybridized modes as exploited in other works [52, 53]. Hybridization of modes adds difficulty in nulling $|g_{\text{eff}}|$, which explains the monotonic increase in the two residuals when $|\Phi_{\text{C2}}|>0.88\Phi_{0}$.
The findings in Figs. 5 imply that flux crosstalk compensation creates a magnetic flux crosstalk-free digital twin of the coupler-mediated CZ gate based on empirical adjustments of the Rabi oscillation model [6, 49, 54]. This digital twin helps identify non-idealities in two-qubit gates such as reduced 2Q Rabi oscillation cycles brought upon by an effective flux transient [6, 31, 49] and unintended hybridization of the qubit and coupler modes. Shaping the flux pulse to account for flux transients [6, 31, 49] while avoiding frequencies that invoke unwanted hybridization of modes [52] could help to match the digital twin with experimental results. We consider the proposed measurement, modeling, and analysis protocol to be suitable for scaling to hundreds of qubits with the aid of high-performance computing [2] (see Sec. SVII of [33] for supporting details).
In this work, we develop the MZLC protocol for characterizing flux crosstalk in a TC architecture. The method exploits residual inductive coupling between several flux-tunable elements to resolve the 0–1 transition frequencies of the probed element. Whereas other flux crosstalk characterization strategies use Rabi spectroscopy  [23, 27] and Ramsey sequences on a single-biased element to characterize the flux crosstalk matrix, relying on the readout fidelity of the qubit/coupler. MZLC has the advantage of characterizing flux crosstalk even in modest readout SNR and an arbitrary idle flux bias range. The method requires knowledge of the dependence of the qubit/coupler spectrum on flux bias, which is routine in the bring-up of superconducting processors, making the MZLC a simple, time-efficient tool for characterizing the flux crosstalk matrix of a transmon-coupler subsystem and optimizing CZ gate operation points. Lastly, compared with the CZ SWAP model, the experimental result shows a nearly identical map with zero crosstalk. Therefore, it provides a way to create a digital twin for the TC system CZ optimization.
Acknowledgment
We acknowledge funding support from the Academia Sinica Grand Challenge project (AS-GCP-112-M01), Grand Challenge Program Seed Grant (AS-GCS-114-M04) National Quantum Initiative (AS-KPQ-111-TQRB), and NSTC (113-2119-M-001-008). C.-T. K. acknowledges the funding support from NSTC 2030 Cross-Generation Young Scholars(112-2628-M-001-004). H.-S. Goan acknowledges support from the National Science and Technology Council (NSTC), Taiwan, under Grants No. NSTC 113-2112-M-002-022-MY3, and No. 114-2119-M-002-018, and from the National Taiwan University under Grants No. NTU-CC-114L8950, and No. NTU-CC114L895004.
C.-D. C. and C.-T. K. conceived the devices and supervised the project with contributions from H.-S. G., S.-Y. H. and C.-H. W.. C.-C. C. and J.-Y. W. designed the device layout with additional inputs from T.-H. L. Y.-C. C. coordinated the device fabrication. Y.-C. C., C.-S. W., C.-D. C., and C.-T. K. provided input on the fabrication methodology. N.-Y. L., L.-C. H. and Y.-S. H. conducted the experiment. M.A.C.A., N.-Y. L., C.-H. M., L.-C. H., and C.-T. K. analyzed the data and wrote the manuscript; all authors discussed the results and contributed to the manuscript.

### v3

@@
 preprint: APL Quantum
-Superconducting qubits are one of the promising types of physical qubits to realize fault-tolerant quantum computing (FTQC), in which quantum computers tackle problems that conventional computers cannot solve on a reasonable time scale [1, 2].
+Superconducting qubits are one of the promising types of physical qubits to realize fault-tolerant quantum computing (FTQC), in which quantum computers tackle problems that conventional computers cannot solve on a reasonable time scale [Arute_2019, Mohseni_2025].
 In operating high-fidelity quantum gates, the tunable coupler plays a crucial role.
-It suppresses spurious coupling and modulates the coupling strength between two qubits [3, 4, 5] via coupler-mediated virtual exchange [4, 1].
-These features enhance single-qubit (1Q) and two-qubit (2Q) gate fidelities in both transmon and Fluxonium architectures  [6, 7] in a superconducting quantum processor.
-Reaching the error correction threshold marks the first step towards FTQC  [1, 8, 9, 10].
+It suppresses spurious coupling and modulates the coupling strength between two qubits [YChen2014, Yan2018, Stehlik_2021] via coupler-mediated virtual exchange [Yan2018, Arute_2019].
+These features enhance single-qubit (1Q) and two-qubit (2Q) gate fidelities in both transmon and fluxonium architectures  [Sung_2021, Ding_2023] in a superconducting quantum processor.
+Reaching the error correction threshold marks the first step towards FTQC  [Arute_2019, Lacroix_2024, Google_2025, Gao_2025].
 To achieve FTQC, an increasing number of qubits and couplers is needed, which leads to more control lines as well as crosstalk between qubits and couplers.
-Crosstalks could originate from unintended microwave [11, 12, 13, 14] and readout drives [15, 16], residual qubit-to-qubit coupling [3, 4, 6, 17], frequency crowding [18, 5, 19, 20], and unwanted flux [21, 22, 23, 24] that leads to imperfection of the 1Q and 2Q gates.
+Crosstalks could originate from unintended microwave [RWang2022, HYan2023, Yang_2024, XYYang2024] and readout drives [Heinsoo_2018, Pitsun_2020], residual qubit-to-qubit coupling [YChen2014, Yan2018, Sung_2021, RLi2024], frequency crowding [Hertzberg2021, Stehlik_2021, EZhang2022, Pappas2024], and unwanted flux [Barends2014, Neill2018, Barrett2023, Dai2024] that leads to imperfection of the 1Q and 2Q gates.
 Flux crosstalk plays a role in a frequency-tunable qubit-coupler system; it obscures the precise frequency detuning required for a quantum gate operation.
 Precise control of the flux pulse avoids unwanted frequency shifts and residual coupling, which is beneficial for quantum gate operation.
-As flux-tunable elements are sources of flux crosstalk in a superconducting processor, the crosstalk matrix shows the spatial effect of the flux source based on the frequency response of other probed elements [23, 25, 26].
+As flux-tunable elements are sources of flux crosstalk in a superconducting processor, the crosstalk matrix shows the spatial effect of the flux source based on the frequency response of other probed elements [Barrett2023, Shi_2023_cal, Xiang_2023].
 The diagonal elements denote the flux source, whereas the off-diagonal elements represent the flux crosstalk.
-Frequency-sensitive techniques like qubit spectroscopy [23, 27], resonator spectroscopy [28, 29, 24], and Ramsey sequences [22, 6, 30] all characterize DC components.
-Phase-sensitive techniques such as Ramsey sequences [28, 27] capture high-frequency components.
-Adding compensating pulses based on the inverse matrix suppresses these crosstalk [21, 22].
-Flux crosstalk compensation, along with flux pulse predistortion [6, 17, 31], enables the prediction of idle frequencies for qubits and couplers in large quantum processors [21, 25, 31].
-In a transmon-coupler (TC) architecture [4, 6, 1, 5], flux control relies on the current flowing through a flux (Z) line grounded near the qubit or coupler.
-As depicted in Fig.
-1(a), four flux lines, colored green and labeled Z, are responsible for frequency control with flux threading through the superconducting quantum interference device (SQUID) loop.
-To characterize flux crosstalk in the TC system, we utilize the tunable two-qubit and two-coupler system, as shown in Fig.
+Frequency-sensitive techniques like qubit spectroscopy [Barrett2023, Kosen_2024], resonator spectroscopy [Abrams_2019, Dai_2021, Dai2024], and Ramsey sequences [Neill2018, Sung_2021, Shi_2023_top] all characterize DC components.
+Phase-sensitive techniques such as Ramsey sequences [Abrams_2019, Kosen_2024] capture high-frequency components.
+Adding compensating pulses based on the inverse matrix suppresses these crosstalk [Barends2014, Neill2018].
+Flux crosstalk compensation, along with flux pulse predistortion [Sung_2021, RLi2024, TMLi2024], enables the prediction of idle frequencies for qubits and couplers in large quantum processors [Barends2014, Shi_2023_cal, TMLi2024].
+The transmon-coupler (TC) architecture [Yan2018, Sung_2021, Arute_2019, Stehlik_2021] offers an experimental testbed for implementing flux crosstalk protocols as its flux control relies on current through flux lines grounded near the qubit or coupler.
+We utilized a tunable two-qubit and two-coupler system to characterize flux crosstalk, as illustrated in Fig.
 1(a).
-Here, the element referring to the flux line $i,k\in\{\textrm{{$Z_{\textrm{Q1}}$},{$Z_{\textrm{Q2}}$}\,$Z_{\textrm{C1}}$},\textrm{$Z_{\textrm{C2}}$}\}$ interacts with an element referring to the SQUID loop of a target qubit or coupler $j\in\{\textrm{Q1},\textrm{Q2},\textrm{C1},\textrm{C2}\}$.
-When the current $I_{i}$ generated by a voltage difference $V_{i}$ across a load impedance of $50\,\Omega$ passes through a target flux line antenna, $I_{i}$ induces two effects that generate flux crosstalk: 1) residual inductive coupling $L_{ij,i\neq j}$ between flux lines and SQUID loops of neighboring qubits and couplers and 2) unwanted fluxes threading through the SQUID structure due to return currents [23, 27].
-Flux crosstalk complicates the precise control of TC system for quantum gate operations.
+The TC archtecture consists of transmon qubits, colored blue and labeled Q, and tunable transmon couplers, colored orange and labeled C; all equipped with flux lines, colored green and labeled Z.
+We study the interaction between the four flux lines $i\in\{\textrm{{$Z_{\textrm{Q1}}$},{$Z_{\textrm{Q2}}$}\,$Z_{\textrm{C1}}$},\textrm{$Z_{\textrm{C2}}$}\}$, and the SQUID loops of the target elements $j\in\{\textrm{Q1},\textrm{Q2},\textrm{C1},\textrm{C2}\}$.
+When the current $I_{i}$ generated by a voltage difference $V_{i}$ across a load impedance of $50\,\Omega$ passes through a target flux line antenna, $I_{i}$ induces two effects that generate flux crosstalk: 1) residual inductive coupling $L_{ij,i\neq j}$ between flux lines and SQUID loops of neighboring qubits and couplers and 2) unwanted fluxes threading through the SQUID structure due to return currents [Barrett2023, Kosen_2024].
+The subsystem is part of a larger quantum processing unit (QPU), an optical micrograph image of which is shown in Fig.
+1(b).
+[…]
 To overcome the crosstalk challenge, we develop a protocol for characterizing flux crosstalk in a compact TC architecture that exploits residual inductive coupling between several flux-tunable elements to resolve the 0–1 transition frequencies of the probed element, named ”Multi-Z-Line Control” (MZLC).
@@
 Color labels are assigned based on the corresponding components in the subsystem’s lumped circuit model shown in (d).
-The protocol for characterizing flux crosstalk is carried out on a subsystem of a quantum processing unit (QPU) containing five qubits and four couplers, as shown in Figs.
-1(b).
-The QPU is fabricated with conventional e-beam lithography, followed by a tilt-angle deposition of aluminum with in situ oxidation to make the circuit elements and Manhattan-style Josephson junctions.
-The MZLC protocol is implemented in a TC subsystem located at the edge of the linear qubit chain, as shown in Figs.
-1(b) and 1(c).
-Wire bonds are placed above the microwave and flux lines to suppress slot line modes [32], and flux crosstalk [27].
-The experiment is conducted in a dilution refrigerator with a base temperature of $10\text{\,}\mathrm{m}\mathrm{K}$, and the microwave signal is supplied by Quantum Machines OPX++\, along with Octave up/down converters (see Sec.
-SII of [33] for additional details).
 […]
-[…]
-Next, we detail the protocol for characterizing the crosstalk of the TC system.
-Figures 2(a) and 2(b) show standard two-tone spectroscopy pulse schedules for qubits and couplers, respectively, with the same color codes as Figs.
+The MZLC protocol requires an initial knowledge of the frequency dependence of the qubits and couplers on flux bias.
+This information is obtained via standard two-tone spectroscopy, using the pulse schedules shown in Figs.
+2(a) and 2(b), respectively, with the same color codes as Figs.
 1.
@@
 2(b).
-ICS exploits the weak capacitive coupling between the coupler electrode and the XY and readout lines of the adjacent qubits to resolve $f_{01,\text{C}}$ [34, 35].
+ICS exploits the weak capacitive coupling between the coupler electrode and the XY and readout lines of the adjacent qubits to resolve $f_{01,\text{C}}$ [Li_2020, collodo_2020].
 In Fig.
@@
 2(c) to 2(f).
-The model is written as  [36, 23]
-$\begin{split}f_{01}(V_{i})&=\left(f_{01,\max}+\frac{E_{C}}{h}\right)\times\\ &\quad\sqrt[4]{d^{2}+\left[1-d^{2}\right]\textrm{cos}^{2}\left(A_{c}\left[V_{i}-V_{\text{ofs,i}}\right]\right)}\\ &\quad-\frac{E_{C}}{h},\end{split}$ (1)
+The model is written as  [Koch2007, Barrett2023]
+$\begin{split}f_{01}(V_{i})&=\left(f_{01,\max}+E_{C}/h\right)\times\\ &\quad\sqrt[4]{d^{2}+\left[1-d^{2}\right]\textrm{cos}^{2}\left(A_{c}\left[V_{i}-V_{\text{ofs,i}}\right]\right)}\\ &\quad-E_{C}/h,\end{split}$ (1)
 where $f_{\text{01,max}}$ is the upper sweet spot frequency, $E_{C}/h$ is the anharmonicity, $d$ is the junction asymmetry, $A_{c}$ is a conversion factor related to the mutual inductance between the $i^{th}$ SQUID loop and the $i^{th}$ flux line, and $V_{\text{ofs,}i}$ is the offset voltage induced by the remnant flux.
@@
 The junction asymmetries for both qubits and couplers are based on the nominal design in Sec.
-SI of [33].
+SI of [SupMat].
 Hence, $A_{c}$, $f_{\text{01,max}}$ and $V_{\text{ofs,}i}$ are extracted when fitting the model with the driven spectra in Figs.
-2(c) and 2(f), expressed in voltages from the analog-to-digital converter in-phase channel (ADC-I) [37].
-These parameters map the idle transmon frequency to its corresponding flux bias  [36], and vice versa after flux crosstalk compensation (see Sec.
-SIII A in [33] for details).
+2(c) and 2(f), expressed in voltages from the analog-to-digital converter in-phase channel (ADC-I) [Sank2025].
+These parameters map the idle transmon frequency to its corresponding flux bias  [Koch2007], and vice versa after flux crosstalk compensation (see Sec.
+SIII A in [SupMat] for details).
 Figure 3: Characterization of flux crosstalk.
@@
 MZLC resolves the behavior of the 0–1 transition of the probe by monitoring changes in its population, which is between $|0\rangle$ and $|1\rangle$ for a weak pulsed XY drive, with different combinations of P-Flux and S-Flux.
-The combined pulses will allow additional sampling of the flux crosstalk effect with better precision, and can ensure the linear correlation between probe and source elements, instead of 1D scans as done in other works [22].
+The combined pulses will allow additional sampling of the flux crosstalk effect with better precision, and can ensure the linear correlation between probe and source elements, instead of 1D scans as done in other works [Neill2018].
 The scheme hinges on detecting flux crosstalk by observing undesirable frequency shifts on the detector induced by S-Flux, which is visually distinct from other sources of unwanted frequency shifts.
@@
 With flux crosstalk, the 0–1 transition appears to vary linearly with the S-Flux sweep.
-The magnitude of the slope captures the total applied flux magnitudes as done previously with other techniques [22].
-Other sources of frequency shift, such as dispersive shifts caused by neighboring flux-tunable elements, and ZZ interaction-induced shift, appear as a nonlinear dependence of the probe’s 0–1 transition [23] at different combinations of simultaneously biased P-Flux and S-Flux.
-It is worth noting that an alternative method is the Ramsey-based method [22, 6, 30], which observes the same sign and magnitude of crosstalk, as shown in Fig.
+The magnitude of the slope captures the total applied flux magnitudes as done previously with other techniques [Neill2018].
+Other sources of frequency shift, such as dispersive shifts caused by neighboring flux-tunable elements, and ZZ interaction-induced shift, appear as a nonlinear dependence of the probe’s 0–1 transition [Barrett2023] at different combinations of simultaneously biased P-Flux and S-Flux.
+It is worth noting that an alternative method is the Ramsey-based method [Neill2018, Sung_2021, Shi_2023_top], which observes the same sign and magnitude of crosstalk, as shown in Fig.
 4(b) and Fig.
-S2 of the supplemental material [33].
+S2 of the supplemental material [SupMat].
 Within the S-flux ranges shown in Figs.
@@
 3(c) and 3(e).
-We attribute these megahertz discrepancies to uncorrected long-time-scale transients of the flux pulses [6, 31, 17] (see Fig.
+We attribute these megahertz discrepancies to uncorrected long-time-scale transients of the flux pulses [Sung_2021, TMLi2024, RLi2024] (see Fig.
 S4 and Sec.
-SIII F of [33] for details).
+SIII F of [SupMat] for details).
 These megahertz discrepancies, if not corrected first with flux predistortion and then flux crosstalk compensation, lead to coherent errors that affect the 1Q gate fidelities of qubits in a multi-qubit processor having simultaneous flux pulse biasing (see Secs.
-SVI and SVII A of [33] for details).
+SVI and SVII A of [SupMat] for details).
 When using the MZLC approach, detuning the probe element away from its upper sweet spot leads to a narrow linewidth, which improves the precision of flux crosstalk measurement, as shown in Fig.
@@
 These linewidths are consistent with the qubit and coupler spectra in Fig.
-2, which are observed in previous works [38, 39, 40, 6, 41].
+2, which are observed in previous works [Schuster_2005, Paik2011, Whittaker2014, Sung_2021, Anferov_2024].
 In fast-flux qubit spectroscopy, whose pulse time is much faster than the relaxation and dephasing time $t\ll T_{1},T_{2}^{*}$, the linewidth of the qubit spectrum is Fourier-limited by the inverse of the pulse duration time.
-However, the XY Rabi drive, which is proportional to the transmon’s dipole matrix element, and consequently the fourth root of the Josephson energy [42], has periodic dependence on the applied flux.
+However, the XY Rabi drive, which is proportional to the transmon’s dipole matrix element, and consequently the fourth root of the Josephson energy [Sung_2022], has periodic dependence on the applied flux.
 The flux-dependent XY Rabi drive modulates the effective qubit linewidth.
-As our detection undergoes repeated ensemble averaging per pixel to resolve the 0–1 transition, low-frequency drifts (like qubit frequency drifts) are averaged out, leading to qubit dephasing [43].
-At a weak XY drive, the observed pulsed spectroscopy lineshape becomes Gaussian-like due to inhomogeneous broadening [44, 43, 45] (see Sec.
-SIV of [33] for details).
+As our detection undergoes repeated ensemble averaging per pixel to resolve the 0–1 transition, low-frequency drifts (like qubit frequency drifts) are averaged out, leading to qubit dephasing [Martinis_2003].
+At a weak XY drive, the observed pulsed spectroscopy lineshape becomes Gaussian-like due to inhomogeneous broadening [Blais_2021, Martinis_2003, Krantz_2019].
 Despite the observed broadening, peak thresholding and curve fitting improve the frequency shift resolution (see Sec.
-SV of [33] for details).
+SV of [SupMat] for details).
 Nonetheless, the sign and amplitude of the crosstalk in Figs.
@@
 Total $\sum_{i\neq k}|X_{ik}|$ 318.4 2.42
-Matrix
-Asymmetry
-$\sum_{i\neq k}$
-$||X_{ik}|-|X_{ki}||$
-143.6 0.94
+Matrix Asymmetry $\sum_{i\neq k}$ $||X_{ik}|-|X_{ki}||$ 143.6 0.94
 We summarize the crosstalk matrices measured using the MZLC and Ramsey sequences in Figs.
@@
 We discuss the efficiency of MZLC and Ramsey schemes in Secs.
-SIII B, and SVII of [33].
+SIII B, and SVII of [SupMat].
 Table 1 shows the statistical properties of $X_{ik}$ before and after flux compensation in Figs.
@@
 First, the average uncompensated crosstalk has the same order as its extreme values (Fig.
-S3 of [33] presents the same results in dB).
-Its total crosstalk is similar to previous flux-tunable two-qubit transmons [46].
+S3 of [SupMat] presents the same results in dB).
+Its total crosstalk is similar to previous flux-tunable two-qubit transmons [Chow2010].
 Its matrix asymmetry exhibits an imbalance likely stemming from the leftward routing layout of our flux lines, producing asymmetric inductive coupling and return current.
-A careful routing of flux lines [27] with the return current flowing directly on the PCB ground [47] could passively reduce the flux crosstalk and matrix asymmetry to a certain extent.
+A careful routing of flux lines [Kosen_2024] with the return current flowing directly on the PCB ground [Niu2024] could passively reduce the flux crosstalk and matrix asymmetry to a certain extent.
 Lastly, these metrics dropped one to two orders of magnitude after flux compensation; residual crosstalk averages near zero, with a standard deviation of a similar scale.
@@
 Panels from (b) to (e) share the same x-axis, representing the applied flux on the coupler.
-In addition to compensating for flux crosstalk, we applied this method to help locate the optimal bias point for conditional phase (CZ) gate operations [48, 6, 17].
+In addition to compensating for flux crosstalk, we applied this method to help locate the optimal bias point for conditional phase (CZ) gate operations [Foxen_2020, Sung_2021, RLi2024].
 In TC architectures, the coupler-mediated CZ gate works by tuning the frequencies of the qubits so that the states $|11\rangle$ and $|20\rangle$ become resonant, resulting in a state swap.
@@
 5(a) (see Sec.
-SIII D of [33] for the relevant CZ SWAP idle frequencies).
+SIII D of [SupMat] for the relevant CZ SWAP idle frequencies).
 The behavior seen in Figs.
 5(b) and 5(c) before and after flux compensation, respectively, are compared with Fig.
-5(d), which is derived from the Rabi oscillation for 2Q gates [41, 49], accounting for the finite rising time of the diabatic flux pulse.
+5(d), which is derived from the Rabi oscillation for 2Q gates [Anferov_2024, hellings_2025], accounting for the finite rising time of the diabatic flux pulse.
 This finite rise time—an effective flux transient representing the combined, rather than individual, components of flux pulse distortions—results in fewer Rabi oscillation cycles and, consequently, a reduced interaction time for two-qubit (2Q) gates (see additional details of the model from Secs.
-SIII D to SIII F of [33]).
-Due to phase errors arising from the uncalibrated pulse distortions in flux biasing [25, 17], Figs.
+SIII D to SIII F of [SupMat]).
+Due to phase errors arising from the uncalibrated pulse distortions in flux biasing [Shi_2023_cal, RLi2024], Figs.
 5(b) and 5(c) may show a small deviation from Fig.
@@
 […]
-The CZ coupling strength in a TC architecture (Q1-C2-Q2) is written as [50, 51]
+The CZ coupling strength in a TC architecture (Q1-C2-Q2) is written as [Sete_2021, Liang_2023_swap]
 $g_{\text{eff}}=\sqrt{2}\left(g_{12}-\frac{g_{1c}g_{2c}}{2}B_{02}\right),$ (3)
@@
 5(e), we use Eqs.
-(3)–(4) with an added coupling strength of |435||435|\,kHz, which could be reduced through better readout fidelity and rigorous flux predistortions [17, 25, 49].
+(3)–(4) with an added coupling strength of |435||435|\,kHz, which could be reduced through better readout fidelity and rigorous flux predistortions [RLi2024, Shi_2023_cal, hellings_2025].
 The inset in Fig.
@@
 The fit cannot explain the sharp peaks at $|g_{\text{eff}}/2\pi|$ at $|\Phi_{\text{C2}}|\approx 0.88\Phi_{0}$ ($f_{\text{C2}}\approx$ $5.94\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$) as the predicted singularities occur beyond the range of the coupler flux displayed in Fig.
-5(e) [51].
-The singularities probably come from the strong coupling between the readout resonator R2 and the coupler C2 as $f_{\text{C2}}\approx f_{\text{R2}}$ (see Table S2 of [33]), which produces hybridized modes as exploited in other works [52, 53].
+5(e) [Liang_2023_swap].
+The singularities probably come from the strong coupling between the readout resonator R2 and the coupler C2 as $f_{\text{C2}}\approx f_{\text{R2}}$ (see Table S2 of [SupMat]), which produces hybridized modes as exploited in other works [chen2024_reset, zhang2025_aswap].
 Hybridization of modes adds difficulty in nulling $|g_{\text{eff}}|$, which explains the monotonic increase in the two residuals when $|\Phi_{\text{C2}}|>0.88\Phi_{0}$.
 The findings in Figs.
-5 imply that flux crosstalk compensation creates a magnetic flux crosstalk-free digital twin of the coupler-mediated CZ gate based on empirical adjustments of the Rabi oscillation model [6, 49, 54].
-This digital twin helps identify non-idealities in two-qubit gates such as reduced 2Q Rabi oscillation cycles brought upon by an effective flux transient [6, 31, 49] and unintended hybridization of the qubit and coupler modes.
-Shaping the flux pulse to account for flux transients [6, 31, 49] while avoiding frequencies that invoke unwanted hybridization of modes [52] could help to match the digital twin with experimental results.
-We consider the proposed measurement, modeling, and analysis protocol to be suitable for scaling to hundreds of qubits with the aid of high-performance computing [2] (see Sec.
-SVII of [33] for supporting details).
+5 imply that flux crosstalk compensation creates a magnetic flux crosstalk-free intuitive digital twin of the coupler-mediated CZ gate.
+This twin is based on empirical adjustments of the Rabi oscillation model [hellings_2025], whose analytical form is validated by carefully calibrated two-qubit state measurements [Krizan_2025] and initially modelled as three-interacting qutrits [Sung_2021].
+This digital twin helps identify non-idealities in two-qubit gates such as reduced 2Q Rabi oscillation cycles brought upon by an effective flux transient [Sung_2021, TMLi2024, hellings_2025] and unintended hybridization of the qubit and coupler modes.
+Shaping the flux pulse to account for flux transients [Sung_2021, TMLi2024, hellings_2025] while avoiding frequencies that invoke unwanted hybridization of modes [chen2024_reset] could help to match the digital twin with experimental results.
+We consider the proposed measurement, modeling, and analysis protocol to be suitable for scaling to hundreds of qubits with the aid of high-performance computing [Mohseni_2025] (see Sec.
+SVII of [SupMat] for supporting details).
 In this work, we develop the MZLC protocol for characterizing flux crosstalk in a TC architecture.
 The method exploits residual inductive coupling between several flux-tunable elements to resolve the 0–1 transition frequencies of the probed element.
-Whereas other flux crosstalk characterization strategies use Rabi spectroscopy  [23, 27] and Ramsey sequences on a single-biased element to characterize the flux crosstalk matrix, relying on the readout fidelity of the qubit/coupler.
+Whereas other flux crosstalk characterization strategies use Rabi spectroscopy  [Barrett2023, Kosen_2024] and Ramsey sequences on a single-biased element to characterize the flux crosstalk matrix, both rely on good readout fidelity of the qubit/coupler.
 MZLC has the advantage of characterizing flux crosstalk even in modest readout SNR and an arbitrary idle flux bias range.
-The method requires knowledge of the dependence of the qubit/coupler spectrum on flux bias, which is routine in the bring-up of superconducting processors, making the MZLC a simple, time-efficient tool for characterizing the flux crosstalk matrix of a transmon-coupler subsystem and optimizing CZ gate operation points.
+The method requires knowledge of the dependence of the qubit (coupler) spectrum on flux bias, which is routine in the bring-up of superconducting processors.
+As discussed in the supplemental material [SupMat], its primary advantage stems from avoiding the prerequisite calibration steps for precise $\pi$ and $\pi/2$ gate operations when calibrating flux crosstalk.
+This combination makes the MZLC a simple, scalable, robust and low-overhead tool for characterizing the flux crosstalk matrix of a transmon-coupler subsystem and optimizing CZ gate operation points.
 Lastly, compared with the CZ SWAP model, the experimental result shows a nearly identical map with zero crosstalk.
 Therefore, it provides a way to create a digital twin for the TC system CZ optimization.
+Supplemental Material
+The supplemental material [SupMat] includes the following: I) details of the QPU fabrication and design; II) experimental details and essential QPU parameters; III) calibration between flux and qubit-coupler interactions; IV) flux-induced inhomogeneous effects on linewidth in MZLC; V) spectroscopic signal-to-noise ratio and frequency-shift uncertainty of MZLC; VI) effect of flux crosstalk compensation on single-qubit gate fidelities; VII) scalability of flux crosstalk characterization and compensation for large-scale superconducting processors.
+Data Availability
+The data that support the most of the findings of this study are openly available in Zenodo at http://doi.org/10.5281/zenodo.17639715.
+Check the latest version for the updated datasets.
+All other data are available from the corresponding authors upon request.
 Acknowledgment
@@
 NTU-CC114L895004.
+Author Declaration
+Conflict of Interest
+The authors have no conflicts to declare.
+Author Contributions
 C.-D.